OpenStudy (anonymous):

If x*x + 2x - 35 = 0, then x = ? If x*x + 2x - 35 = 0, then x = ? @Mathematics

OpenStudy (anonymous):

x = -7 or, 5

OpenStudy (anonymous):

or 5?

OpenStudy (anonymous):

(x+7)(x-5) = 0 either, x+7 = 0 or, x-5 = 0 x = -7 x = 5

OpenStudy (ash2326):

yeah both of the values satisfy the equation , if we have a condition on x then we can rule out anyone of them . say x should be positive then only solution is 5 and likewise

OpenStudy (anonymous):

ok thanksss ;]

OpenStudy (anonymous):

x =5 or -7